# Maximum entropy based testing in network models: ERGMs and constrained optimization

### v1

@@ Abstract
Stochastic network models play a central role across a wide range of scientific disciplines, and questions of statistical inference arise naturally in this context. In this paper we investigate goodness-of-fit and two-sample testing procedures for statistical networks based on the principle of maximum entropy (MaxEnt). Our approach formulates a constrained entropy-maximization problem on the space of networks, subject to prescribed structural constraints. The resulting test statistics are defined through the Lagrange multipliers associated with the constrained optimization problem, which, to our knowledge, is novel in the statistical networks literature.
We establish consistency in the classical regime where the number of vertices is fixed. We then consider asymptotic regimes in which the graph size grows with the sample size, developing tests for both dense and sparse settings. In the dense case, we analyze exponential random graph models (ERGM) (including the Erdös-Rènyi models), while in the sparse regime our theory applies to Erdös-Rènyi graphs.
Our analysis leverages recent advances in nonlinear large deviation theory for random graphs. We further show that the proposed Lagrange-multiplier framework connects naturally to classical score tests for constrained maximum likelihood estimation. The results provide a unified entropy-based framework for network model assessment across diverse growth regimes.
@@ Introduction
In the context of goodness-of-fit testing, various approaches have been developed for different random graph models. For instance, [50] and [45] proposed motif-based tests for bipartite network models and kernel-based tests for general random graphs, respectively. For inhomogeneous random graph models, a degree-based test has been introduced in [51], while [22] develops a test based on the adjacency matrix of the random graph. In the context of the stochastic block model, [43] also develops an adjacency matrix-based test. However, unlike [22], which considers multiple independent samples, [43] focuses on a test based on a single sample. Several other methods have also been explored. For example, [68] studies goodness-of-fit testing for exponential random graph models (ERGMs) and introduces a test based on kernel Stein discrepancies. Additionally, [12] presents a framework for testing the homogeneity of a random graph using graph functionals, which generalizes the subgraph count approach.
@@
Although our analysis is centered on strictly balanced graph counts, which includes important motifs such as cliques and cycles (see the notation section for a definition of strictly balanced graphs), the arguments can, in principle, be extended to general motif counts. However, we focus on a strictly balanced graph to maintain clarity of exposition. Moreover, while our study is restricted to the Erdös-Rènyi model (both sparse and dense) and ERGMs (in the dense regime), the underlying principle of entropy maximization is quite general and may be adapted to other random graph models. However, even in these cases, the analysis of the Lagrange multiplier is technically involved, and we leave the extensions to the more general models for future work.
@@
In Section 2 we describe the ERGM model and the general hypothesis testing problem. We also describe our testing procedure based on the Lagrange multiplier which is in turn obtained from the empirical entropy maximization under constraints on motif counts. The analysis in the dense regime uses the theory of graphons and hence we also include a brief introduction to graphons in Section 2. Section 3 contains main results of our paper while Section 4 is devoted to a short discussion on the key ideas used in the proofs. Finally, we conclude in Section 5 with a discussion on possible generalizations of our results. The proofs of the main theorems are given in Appendix A while the technical lemmas underpinning our main arguments are provided in the Appendix C.
@@
    Finally, our proposed tests are based on a single Lagrange multiplier. In contrast, the framework of [2] suggests the possibility of tests based on a vector of multipliers arising from hypotheses involving multiple parameters. While this generalization appears promising, it lies beyond the scope of the present work and will be explored in future research.
@@
ERGMs grew from exponential-family models in social network analysis [29, 38, 65], and since then this model has been widely studied (cf. [52, 54, 15, 55, 56, 71, 58, 24, 47, 27, 28, 7, 57, 9, 30, winstein2025concentration, 66] for a partial list). While the literature on this topic is extensive (and the above list is by no means exhaustive), we note that hypothesis testing within the ERGM framework has been relatively underexplored. In particular, while a few recent works such as [68] and [57] develop tests based on Stein discrepancies and [69] proposes tests based on sums of degrees, the wider literature on the topic is rather sparse.
@@
In this regime, ERGMs asymptotically resemble an Erdös-Rènyi $G(N,p)$ and is amenable to analysis. On the other hand, if the parameter $\beta$ lies outside this region (in other words, it lies in the supercritical regime), the graphs exhibit clustering and metastability, and are much harder to analyze (see [winstein2025concentration, 66] for details). We will restrict ourselves to the subcritical regime of a ferromagnetic ERGM and will use the tools from [15] to analyze the Lagrange multiplier and thus develop a test.
@@
Understanding the behavior of the log normalizing constant is a crucial aspect in our analysis of the Lagrange multiplier. Although details are given in Section 4, we note here that the solution of (2.6) can be seen as a critical point of the function $\sum_{i=1}^{n}e^{-{\lambda}h_{n,i}}$. It can be shown that the expected value of the same function is exactly the normalizing constant $Z_{N}(\beta)$ for a certain ERGM and hence (2.20) features heavily in the analysis of our solution $\hat{\lambda}_{n}$. However, the above argument is at the expectation level and the same argument cannot be easily generalized to analyze the empirical version, and we need more nuanced tools.
@@
In this section, we present the main results of our paper. The main goal will be to to construct the goodness-of-fit and the two sample tests by analyzing the asymptotic behaviour of the Lagrange optimizer $\hat{\lambda}_{n}$.
@@
Proof Outline
@@
    $\mathbb{E}[(H_{m}-h_{0})e^{-\lambda(H_{m}-h_{0})}]=\mathbb{E}[(Z_{\mu}-h_{0})e^{-\lambda(Z_{\mu}-h_{0})}](1\pm o(1)).$ (4.1)
@@
    Using standard arguments from empirical process theory and (4.1), we show that
@@
    $\sqrt{n}(\hat{\lambda}_{n}-\lambda^{\circ})=\frac{-\sqrt{n}(\Psi_{n}(\lambda^{\circ})-\mathbb{E}[\Psi_{n}(\lambda^{\circ})])}{\dot{\Psi}_{n}(\lambda^{\circ})+\frac{1}{2}(\hat{\lambda}_{n}-\lambda^{\circ})\ddot{\Psi}_{n}(\tilde{\lambda}_{n})}+o_{p}(1)$ (4.2)
@@
    Then the consistency of $\hat{\lambda}_{n}$ ensures that the denominator in the RHS of (4.2) converges in probability:
@@
    $\frac{1}{m_{n}^{2}}\log\left(\sum_{G}\frac{N(G)}{n}e^{-\frac{\lambda}{m_{n}^{v(H)-2}}(\mathcal{T}(H,G)-h_{0})}\right)=\frac{\lambda h_{0}}{m_{n}^{v(H)}}+\frac{1}{m_{n}^{2}}\log\left(\sum_{x\in\{0,1\}^{\binom{m_{n}}{2}}}e^{f(x)}\right),$ (4.4)
@@
    $-\sup_{\tilde{h}\in\tilde{W}}\left(\sum_{k=1}^{K}\beta_{k}t(T_{k},h)-\frac{1}{2}I(h)\right)+o(1).$ (4.5)
    Combining (4.4) with (4.5) we obtain
@@
    $m_{n}^{2}\,\hat{\lambda}_{n}=\frac{\frac{1}{n}\displaystyle\sum_{i=1}^{n}\frac{h_{n,i}}{m_{n}^{v(H)}}}{\frac{1}{n}\displaystyle\sum\limits_{i=1}^{n}\frac{h_{n,i}^{2}}{m_{n}^{2v(H)}}e^{-\frac{\tilde{\lambda}_{n}}{m_{n}^{v(H)-2}}h_{n,i}}}=\frac{\frac{1}{\sqrt{n}}\displaystyle\sum_{i=1}^{n}\frac{h_{n,i}}{m_{n}^{v(H)-1}}}{\frac{1}{n}\displaystyle\sum\limits_{i=1}^{n}\frac{h_{n,i}^{2}}{m_{n}^{2v(H)}}e^{-\frac{\tilde{\lambda}_{n}}{m_{n}^{v(H)-2}}h_{n,i}}}\frac{1}{\sqrt{n}{m_{n}}}.$ (4.6)
@@
    $\frac{m_{n}}{\sqrt[4]{n}}Z_{n}\xrightarrow{P}0.$ (4.7)
@@
    $V_{n}:=\sqrt[4]{n}\times m_{n}\times\frac{1}{n}\sum_{i=1}^{n}\frac{h_{n,i}^{2}}{m_{n}^{2v(H)}}e^{-\frac{\tilde{\lambda}_{n}}{m_{n}^{v(H)-2}}h_{n,i}}=\Omega_{P}\left(\frac{1}{m_{n}}\right).$ (4.8)
@@
    Finally, we plug-in (4.7) and (4.8) in (4.6) to conclude that $m_{n}^{2}\hat{\lambda}_{n}\xrightarrow{P}0$.
@@
    $\frac{\sum\limits_{i=1}^{n}\frac{-h_{n,i}}{m_{n}^{v(H)}}e^{-\frac{\lambda^{\circ}}{m_{n}^{v(H)-2}}h_{n,i}}}{\sum\limits_{i=1}^{n}e^{-\frac{\lambda^{\circ}}{m_{n}^{v(H)-2}}h_{n,i}}}\xrightarrow{P}\frac{d}{d\lambda}\mathfrak{g}(\lambda)\bigg|_{\lambda=\lambda^{\circ}}=\frac{p_{0}^{e(H)}}{|Aut(H)|}-\frac{(u^{*})^{e(H)}}{|Aut(H)|}$ (4.9)
@@
    $\frac{\sum\limits_{i=1}^{n}\frac{h_{n,i}^{2}}{m_{n}^{2v(H)}}e^{-\frac{\lambda^{\circ}}{m_{n}^{v(H)-2}}h_{n,i}}}{\sum\limits_{i=1}^{n}e^{-\frac{\lambda^{\circ}}{m_{n}^{v(H)-2}}h_{n,i}}}\xrightarrow{P}\left(\frac{(u^{*})^{e(H)}}{|Aut(H)|}-\frac{p_{0}^{e(H)}}{|Aut(H)|}\right)^{2}$ (4.10)
@@
    $m_{n}^{2}(\hat{\lambda}_{n}-\lambda^{\circ})\xrightarrow{P}\frac{\frac{(u^{*})^{e(H)}}{|Aut(H)|}-\frac{p_{0}^{e(H)}}{|Aut(H)|}}{\left(\frac{(u^{*})^{e(H)}}{|Aut(H)|}-\frac{p_{0}^{e(H)}}{|Aut(H)|}\right)^{2}}=\frac{1}{\frac{(u^{*})^{e(H)}}{|Aut(H)|}-\frac{p_{0}^{e(H)}}{|Aut(H)|}}.$ (4.11)
@@
Discussion
Acknowledgements
Appendix A Proofs
Before we discuss the proofs of the main theorems, we recall some graph theoretic concepts from the Notations subsection.
Homomorphisms and injective embeddings
Labeled and unlabeled copy counts (of the motif $H$ in $G$)
Finally, $(N)_{k}=N^{k}\bigl(1+O_{k}(1/N)\bigr)$, so
@@ Appendix A Proofs
exists and is unique by Lemma D.2 and D.1. Then by Hurwitz’s theorem, $\hat{\lambda}_{n}\rightarrow\lambda^{\circ}$ almost surely, where $\lambda^{\circ}$ is the unique zero of $\sum ae^{-\lambda a}\mathbb{P}_{\mathcal{G}}(h=a)$. ∎
@@
$\mathbb{E}[(H_{m}-h_{0})e^{-\lambda(H_{m}-h_{0})}]=\mathbb{E}[(Z_{\mu}-h_{0})e^{-\lambda(Z_{\mu}-h_{0})}](1\pm o(1))$ (A.1)
Indeed if (A.1) is true, the two functions will be of the same sign for large enough $m$ implying that the LHS is positive at $\lambda^{\circ}-\epsilon$ and negative at $\lambda^{\circ}+\epsilon$ . This will imply that $|\lambda_{m}-\lambda^{\circ}|<\epsilon$. To show (A.1), first note that for $\lambda>0$, the function $x\mapsto(x-h_{0})e^{-\lambda(x-h_{0})}$ is bounded. Hence, by Poisson convergence of $H$ counts in the regime $mp^{k(H)}\rightarrow c>0$, we have that
@@
Noting that in this regime, we also have $\mathbb{E}[(Z_{\mu}-h_{0})e^{-\lambda(Z_{\mu}-h_{0})}]=\Theta(1)$ for fixed $\lambda$, (A.1) holds.
@@
$\sqrt{n}(\hat{\lambda}_{n}-\lambda^{\circ})=\frac{-\sqrt{n}(\Psi_{n}(\lambda^{\circ})-\mathbb{E}[\Psi_{n}(\lambda^{\circ})])}{\dot{\Psi}_{n}(\lambda^{\circ})+\frac{1}{2}(\hat{\lambda}_{n}-\lambda^{\circ})\ddot{\Psi}_{n}(\tilde{\lambda}_{n})}+\frac{-\sqrt{n}\mathbb{E}[\Psi_{n}(\lambda^{\circ})]}{\dot{\Psi}_{n}(\lambda^{\circ})+\frac{1}{2}(\hat{\lambda}_{n}-\lambda^{\circ})\ddot{\Psi}_{n}(\tilde{\lambda}_{n})}.$ (A.2)
@@
Then the consistency of $\hat{\lambda}_{n}$ ensures that the denominators in (A.2) converge almost surely to
$\mathbb{E}\left[-(Z_{\mu}-h_{0})^{2}e^{-\lambda^{\circ}(Z_{\mu}-h_{0})}\right]$ . For the numerators, observe that by arguing as we did for expectation, the variance of $h_{n,1}e^{-\lambda^{\circ}h_{n,1}}$ converges to that of $(Z_{\mu}-h_{0})^{2}e^{-\lambda(Z_{\mu}-h_{0})}$. Hence, the first numerator coverges, by Lemma 3.1, to
@@
$m_{n}^{2}\,\hat{\lambda}_{n}=\frac{\frac{1}{n}\displaystyle\sum_{i=1}^{n}\frac{h_{n,i}}{m_{n}^{v(H)}}}{\frac{1}{n}\displaystyle\sum\limits_{i=1}^{n}\frac{h_{n,i}^{2}}{m_{n}^{2v(H)}}e^{-\frac{\tilde{\lambda}_{n}}{m_{n}^{v(H)-2}}h_{n,i}}}=\frac{\frac{1}{\sqrt{n}}\displaystyle\sum_{i=1}^{n}\frac{h_{n,i}}{m_{n}^{v(H)-1}}}{\frac{1}{n}\displaystyle\sum\limits_{i=1}^{n}\frac{h_{n,i}^{2}}{m_{n}^{2v(H)}}e^{-\frac{\tilde{\lambda}_{n}}{m_{n}^{v(H)-2}}h_{n,i}}}\frac{1}{\sqrt{n}{m_{n}}}.$ (A.4)
@@
$\frac{m_{n}}{\sqrt[4]{n}}Z_{n}\xrightarrow{P}0.$ (A.5)
@@
$V_{n}:=\sqrt[4]{n}\times m_{n}\times\frac{1}{n}\sum_{i=1}^{n}\frac{h_{n,i}^{2}}{m_{n}^{2v(H)}}e^{-\frac{\tilde{\lambda}_{n}}{m_{n}^{v(H)-2}}h_{n,i}}=\Omega_{P}\left(\frac{1}{m_{n}}\right).$ (A.6)
@@
Finally, we plug-in (A.5) and (A.6) in (A.4) to conclude that $m_{n}^{2}\hat{\lambda}_{n}\xrightarrow{P}0$.
@@
$\frac{\sum\limits_{i=1}^{n}\frac{-h_{n,i}}{m_{n}^{v(H)}}e^{-\frac{\lambda^{\circ}}{m_{n}^{v(H)-2}}h_{n,i}}}{\sum\limits_{i=1}^{n}e^{-\frac{\lambda^{\circ}}{m_{n}^{v(H)-2}}h_{n,i}}}\xrightarrow{P}\frac{d}{d\lambda}\mathfrak{g}(\lambda)\bigg|_{\lambda=\lambda^{\circ}}=\frac{p_{0}^{e(H)}}{|Aut(H)|}-\frac{(u^{*})^{e(H)}}{|Aut(H)|}$ (A.7)
@@
$\frac{\sum\limits_{i=1}^{n}\frac{h_{n,i}^{2}}{m_{n}^{2v(H)}}e^{-\frac{\lambda^{\circ}}{m_{n}^{v(H)-2}}h_{n,i}}}{\sum\limits_{i=1}^{n}e^{-\frac{\lambda^{\circ}}{m_{n}^{v(H)-2}}h_{n,i}}}\xrightarrow{P}\left(\frac{(u^{*})^{e(H)}}{|Aut(H)|}-\frac{p_{0}^{e(H)}}{|Aut(H)|}\right)^{2}$ (A.8)

### v2

@@ Abstract
Stochastic network models play a central role across a broad range of scientific disciplines, and questions of statistical inference arise naturally in this context. Maximum entropy (abbrv. MaxEnt) principles are well-regarded as a standard approach to such inferential questions on network data.
In this paper, we investigate goodness-of-fit and two-sample testing procedures for statistical networks grounded in the principle of maximum entropy. Our approach is based on considering a constrained entropy-maximization problem on the space of networks, subject to a prescribed set of structural constraints (informed, for instance, by the goodness-of-fit setup). Our test statistics are based on the Lagrange multipliers associated with such constrained optimization problems, which to our knowledge is a novelty in the statistical networks literature.
We first establish consistency of the proposed tests in the classical setting where the number of vertices is fixed. We then turn to regimes of greater contemporary interest, wherein the size of the graph grows commensurately with the sample size, and develop tests in both the sparse and dense settings. In the dense setting, we analyze graphs generated from exponential random graph models (abbrv. ERGM) as well as Erdös-Rènyi models, while in the sparse regime our theory applies to Erdös-Rènyi graphs. The scope of our results are commensurate with the development of the theory of large scale network asymptotics, varying in robustness from dense to sparse regimes.
Our techniques strongly interface with the burgeoning theory of non-linear large deviations in the space of random graphs, an approach which we believe would be of much wider applicability in statistical network problems. In a different direction, our inferential approach based on Lagrange multipliers turns out to connect to classical score tests for constrained MLE in real-valued data that are widely popular in econometrics. Our results provide a unified entropy-based inferential framework for network model assessment across a wide range of network growth regimes. More broadly, it points to a general overarching strategy whereby goodness-of-fit testing can be framed as a constrained optimization problem and the associated Lagrange multiplier can be used for devising the test.
@@ Introduction
In the context of goodness-of-fit testing, various approaches have been developed for different random graph models. For instance, [56] and [51] proposed motif-based tests for bipartite network models and kernel-based tests for general random graphs, respectively. For inhomogeneous random graph models, a degree-based test has been introduced in [57], while [25] develops a test based on the adjacency matrix of the random graph. In the context of the stochastic block model, [49] also develops an adjacency matrix-based test. However, unlike [25], which considers multiple independent samples, [49] focuses on a test based on a single sample. Several other methods have also been explored. For example, [74] studies goodness-of-fit testing for exponential random graph models (ERGMs) and introduces a test based on kernel Stein discrepancies. Additionally, [14] presents a framework for testing the homogeneity of a random graph using graph functionals, which generalizes the subgraph count approach. We also refer the reader to [10, 15, 9, 37, 34, 36, 35] for related inference problems in other statistical network models.
@@
Although our analysis is centered on strictly balanced graph counts, which include important motifs such as cliques and cycles (see the notation section for a definition of strictly balanced graphs), the arguments can, in principle, be extended to general motif counts. However, we focus on a strictly balanced graph to maintain clarity of exposition. Moreover, while our study is restricted to the Erdös-Rènyi model (both sparse and dense) and ERGMs (in the dense regime), the underlying principle of entropy maximization is quite general and may be adapted to other random graph models. However, even in these cases, the analysis of the Lagrange multiplier is technically involved, and we leave the extensions to the more general models for future work.
@@
In Section 2 we describe the ERGM model and the general hypothesis testing problem. We also describe our testing procedure based on the Lagrange multiplier which is in turn obtained from the empirical entropy maximization under constraints on motif counts. The analysis in the dense regime uses the theory of graphons and hence we also include a brief introduction to graphons in Section 2. Section 3 contains main results of our paper, while the proofs are deferred to the Appendix. Finally, we conclude in Section 4 with a discussion on possible generalizations of our results.
@@
    We construct our test based on a scalar Lagrange multiplier. In contrast, the framework of [2] suggests the possibility of tests based on a vector of multipliers arising from hypotheses involving multiple parameters. While this generalization appears promising, it lies beyond the scope of the present work and will be explored in future research.
  4. 4.
    Finally, in conjunction with [38] (see also the classical works [2, 67]), this article shows that the Lagrange multiplier obtained from maximizing empirical entropy (or empirical likelihood in the case of [38]) can be used effectively for statistical inference in complex models. More broadly, it points to a general overarching strategy whereby goodness-of-fit testing can be framed as a constrained optimization problem and the associated Lagrange multiplier can be used for devising the test.
@@
ERGMs grew from exponential-family models in social network analysis [32, 44, 71], and since then this model has been widely studied (cf. [58, 60, 18, 61, 62, 77, 64, 27, 53, 30, 31, 7, 63, 11, 33, winstein2025concentration, 72] for a partial list). While the literature on this topic is extensive (and the above list is by no means exhaustive), we note that hypothesis testing within the ERGM framework has been relatively underexplored. In particular, while a few recent works such as [74] and [63] develop tests based on Stein discrepancies and [75] proposes tests based on sums of degrees, the wider literature on the topic is rather sparse.
@@
In this regime, ERGMs asymptotically resemble an Erdös-Rènyi $G(N,p)$ and is amenable to analysis. On the other hand, if the parameter $\beta$ lies outside this region (in other words, it lies in the supercritical regime), the graphs exhibit clustering and metastability, and are much harder to analyze (see [winstein2025concentration, 72] for details). We will restrict ourselves to the subcritical regime of a ferromagnetic ERGM and will use the tools from [18] to analyze the Lagrange multiplier and thus develop a test.
@@
Understanding the behavior of the log normalizing constant is a crucial aspect in our analysis of the Lagrange multiplier. Although details are given in Appendix B, we note here that the solution of (2.6) can be seen as a critical point of the function $\sum_{i=1}^{n}e^{-{\lambda}h_{n,i}}$. It can be shown that the expected value of the same function is exactly the normalizing constant $Z_{N}(\beta)$ for a certain ERGM and hence (2.20) features heavily in the analysis of our solution $\hat{\lambda}_{n}$. However, the above argument is at the expectation level and the same argument cannot be easily generalized to analyze the empirical version, and we need more nuanced tools.
@@
In this section, we present the main results of our paper. The main goal will be to construct the goodness-of-fit and the two sample tests by analyzing the asymptotic behaviour of the Lagrange optimizer $\hat{\lambda}_{n}$.
@@
Discussion
Acknowledgements
Appendix A Notational Preliminaries
Before we discuss the proofs of the main theorems, we recall some graph theoretic concepts from the Notations subsection and elaborate them in this section.
Homomorphisms and injective embeddings
Labeled and unlabeled copy counts (of the motif $H$ in $G$)
Finally, $(N)_{k}=N^{k}\bigl(1+O_{k}(1/N)\bigr)$, so
Appendix B Proof Outline
@@
    $\mathbb{E}[(H_{m}-h_{0})e^{-\lambda(H_{m}-h_{0})}]=\mathbb{E}[(Z_{\mu}-h_{0})e^{-\lambda(Z_{\mu}-h_{0})}](1\pm o(1)).$ (B.1)
@@
    Using standard arguments from empirical process theory and (B.1), we show that
@@
    $\sqrt{n}(\hat{\lambda}_{n}-\lambda^{\circ})=\frac{-\sqrt{n}(\Psi_{n}(\lambda^{\circ})-\mathbb{E}[\Psi_{n}(\lambda^{\circ})])}{\dot{\Psi}_{n}(\lambda^{\circ})+\frac{1}{2}(\hat{\lambda}_{n}-\lambda^{\circ})\ddot{\Psi}_{n}(\tilde{\lambda}_{n})}+o_{p}(1)$ (B.2)
@@
    Then the consistency of $\hat{\lambda}_{n}$ ensures that the denominator in the RHS of (B.2) converges in probability:
@@
    $\frac{1}{m_{n}^{2}}\log\left(\sum_{G}\frac{N(G)}{n}e^{-\frac{\lambda}{m_{n}^{v(H)-2}}(\mathcal{T}(H,G)-h_{0})}\right)=\frac{\lambda h_{0}}{m_{n}^{v(H)}}+\frac{1}{m_{n}^{2}}\log\left(\sum_{x\in\{0,1\}^{\binom{m_{n}}{2}}}e^{f(x)}\right),$ (B.4)
@@
    $-\sup_{\tilde{h}\in\tilde{W}}\left(\sum_{k=1}^{K}\beta_{k}t(T_{k},h)-\frac{1}{2}I(h)\right)+o(1).$ (B.5)
    Combining (B.4) with (B.5) we obtain
@@
    $m_{n}^{2}\,\hat{\lambda}_{n}=\frac{\frac{1}{n}\displaystyle\sum_{i=1}^{n}\frac{h_{n,i}}{m_{n}^{v(H)}}}{\frac{1}{n}\displaystyle\sum\limits_{i=1}^{n}\frac{h_{n,i}^{2}}{m_{n}^{2v(H)}}e^{-\frac{\tilde{\lambda}_{n}}{m_{n}^{v(H)-2}}h_{n,i}}}=\frac{\frac{1}{\sqrt{n}}\displaystyle\sum_{i=1}^{n}\frac{h_{n,i}}{m_{n}^{v(H)-1}}}{\frac{1}{n}\displaystyle\sum\limits_{i=1}^{n}\frac{h_{n,i}^{2}}{m_{n}^{2v(H)}}e^{-\frac{\tilde{\lambda}_{n}}{m_{n}^{v(H)-2}}h_{n,i}}}\frac{1}{\sqrt{n}{m_{n}}}.$ (B.6)
@@
    $\frac{m_{n}}{\sqrt[4]{n}}Z_{n}\xrightarrow{P}0.$ (B.7)
@@
    $V_{n}:=\sqrt[4]{n}\times m_{n}\times\frac{1}{n}\sum_{i=1}^{n}\frac{h_{n,i}^{2}}{m_{n}^{2v(H)}}e^{-\frac{\tilde{\lambda}_{n}}{m_{n}^{v(H)-2}}h_{n,i}}=\Omega_{P}\left(\frac{1}{m_{n}}\right).$ (B.8)
@@
    Finally, we plug-in (B.7) and (B.8) in (B.6) to conclude that $m_{n}^{2}\hat{\lambda}_{n}\xrightarrow{P}0$.
@@
    $\frac{\sum\limits_{i=1}^{n}\frac{-h_{n,i}}{m_{n}^{v(H)}}e^{-\frac{\lambda^{\circ}}{m_{n}^{v(H)-2}}h_{n,i}}}{\sum\limits_{i=1}^{n}e^{-\frac{\lambda^{\circ}}{m_{n}^{v(H)-2}}h_{n,i}}}\xrightarrow{P}\frac{d}{d\lambda}\mathfrak{g}(\lambda)\bigg|_{\lambda=\lambda^{\circ}}=\frac{p_{0}^{e(H)}}{|Aut(H)|}-\frac{(u^{*})^{e(H)}}{|Aut(H)|}$ (B.9)
@@
    $\frac{\sum\limits_{i=1}^{n}\frac{h_{n,i}^{2}}{m_{n}^{2v(H)}}e^{-\frac{\lambda^{\circ}}{m_{n}^{v(H)-2}}h_{n,i}}}{\sum\limits_{i=1}^{n}e^{-\frac{\lambda^{\circ}}{m_{n}^{v(H)-2}}h_{n,i}}}\xrightarrow{P}\left(\frac{(u^{*})^{e(H)}}{|Aut(H)|}-\frac{p_{0}^{e(H)}}{|Aut(H)|}\right)^{2}$ (B.10)
@@
    $m_{n}^{2}(\hat{\lambda}_{n}-\lambda^{\circ})\xrightarrow{P}\frac{\frac{(u^{*})^{e(H)}}{|Aut(H)|}-\frac{p_{0}^{e(H)}}{|Aut(H)|}}{\left(\frac{(u^{*})^{e(H)}}{|Aut(H)|}-\frac{p_{0}^{e(H)}}{|Aut(H)|}\right)^{2}}=\frac{1}{\frac{(u^{*})^{e(H)}}{|Aut(H)|}-\frac{p_{0}^{e(H)}}{|Aut(H)|}}.$ (B.11)
@@
Appendix C Proofs of the main Theorems
@@
exists and is unique by Lemma F.2 and F.1. Then by Hurwitz’s theorem, $\hat{\lambda}_{n}\rightarrow\lambda^{\circ}$ almost surely, where $\lambda^{\circ}$ is the unique zero of $\sum ae^{-\lambda a}\mathbb{P}_{\mathcal{G}}(h=a)$. ∎
@@
$\mathbb{E}[(H_{m}-h_{0})e^{-\lambda(H_{m}-h_{0})}]=\mathbb{E}[(Z_{\mu}-h_{0})e^{-\lambda(Z_{\mu}-h_{0})}](1\pm o(1))$ (C.1)
Indeed if (C.1) is true, the two functions will be of the same sign for large enough $m$ implying that the LHS is positive at $\lambda^{\circ}-\epsilon$ and negative at $\lambda^{\circ}+\epsilon$ . This will imply that $|\lambda_{m}-\lambda^{\circ}|<\epsilon$. To show (C.1), first note that for $\lambda>0$, the function $x\mapsto(x-h_{0})e^{-\lambda(x-h_{0})}$ is bounded. Hence, by Poisson convergence of $H$ counts in the regime $mp^{k(H)}\rightarrow c>0$, we have that
@@
Noting that in this regime, we also have $\mathbb{E}[(Z_{\mu}-h_{0})e^{-\lambda(Z_{\mu}-h_{0})}]=\Theta(1)$ for fixed $\lambda$, (C.1) holds.
@@
$\sqrt{n}(\hat{\lambda}_{n}-\lambda^{\circ})=\frac{-\sqrt{n}(\Psi_{n}(\lambda^{\circ})-\mathbb{E}[\Psi_{n}(\lambda^{\circ})])}{\dot{\Psi}_{n}(\lambda^{\circ})+\frac{1}{2}(\hat{\lambda}_{n}-\lambda^{\circ})\ddot{\Psi}_{n}(\tilde{\lambda}_{n})}+\frac{-\sqrt{n}\mathbb{E}[\Psi_{n}(\lambda^{\circ})]}{\dot{\Psi}_{n}(\lambda^{\circ})+\frac{1}{2}(\hat{\lambda}_{n}-\lambda^{\circ})\ddot{\Psi}_{n}(\tilde{\lambda}_{n})}.$ (C.2)
@@
Then the consistency of $\hat{\lambda}_{n}$ ensures that the denominators in (C.2) converge almost surely to
$\mathbb{E}\left[-(Z_{\mu}-h_{0})^{2}e^{-\lambda^{\circ}(Z_{\mu}-h_{0})}\right]$ . For the numerators, observe that by arguing as we did for expectation, the variance of $h_{n,1}e^{-\lambda^{\circ}h_{n,1}}$ converges to that of $(Z_{\mu}-h_{0})^{2}e^{-\lambda(Z_{\mu}-h_{0})}$. Hence, the first numerator converges, by Lemma 3.1, to
@@
$m_{n}^{2}\,\hat{\lambda}_{n}=\frac{\frac{1}{n}\displaystyle\sum_{i=1}^{n}\frac{h_{n,i}}{m_{n}^{v(H)}}}{\frac{1}{n}\displaystyle\sum\limits_{i=1}^{n}\frac{h_{n,i}^{2}}{m_{n}^{2v(H)}}e^{-\frac{\tilde{\lambda}_{n}}{m_{n}^{v(H)-2}}h_{n,i}}}=\frac{\frac{1}{\sqrt{n}}\displaystyle\sum_{i=1}^{n}\frac{h_{n,i}}{m_{n}^{v(H)-1}}}{\frac{1}{n}\displaystyle\sum\limits_{i=1}^{n}\frac{h_{n,i}^{2}}{m_{n}^{2v(H)}}e^{-\frac{\tilde{\lambda}_{n}}{m_{n}^{v(H)-2}}h_{n,i}}}\frac{1}{\sqrt{n}{m_{n}}}.$ (C.4)
@@
$\frac{m_{n}}{\sqrt[4]{n}}Z_{n}\xrightarrow{P}0.$ (C.5)
@@
$V_{n}:=\sqrt[4]{n}\times m_{n}\times\frac{1}{n}\sum_{i=1}^{n}\frac{h_{n,i}^{2}}{m_{n}^{2v(H)}}e^{-\frac{\tilde{\lambda}_{n}}{m_{n}^{v(H)-2}}h_{n,i}}=\Omega_{P}\left(\frac{1}{m_{n}}\right).$ (C.6)
@@
Finally, we plug-in (C.5) and (C.6) in (C.4) to conclude that $m_{n}^{2}\hat{\lambda}_{n}\xrightarrow{P}0$.
@@
$\frac{\sum\limits_{i=1}^{n}\frac{-h_{n,i}}{m_{n}^{v(H)}}e^{-\frac{\lambda^{\circ}}{m_{n}^{v(H)-2}}h_{n,i}}}{\sum\limits_{i=1}^{n}e^{-\frac{\lambda^{\circ}}{m_{n}^{v(H)-2}}h_{n,i}}}\xrightarrow{P}\frac{d}{d\lambda}\mathfrak{g}(\lambda)\bigg|_{\lambda=\lambda^{\circ}}=\frac{p_{0}^{e(H)}}{|Aut(H)|}-\frac{(u^{*})^{e(H)}}{|Aut(H)|}$ (C.7)
@@
$\frac{\sum\limits_{i=1}^{n}\frac{h_{n,i}^{2}}{m_{n}^{2v(H)}}e^{-\frac{\lambda^{\circ}}{m_{n}^{v(H)-2}}h_{n,i}}}{\sum\limits_{i=1}^{n}e^{-\frac{\lambda^{\circ}}{m_{n}^{v(H)-2}}h_{n,i}}}\xrightarrow{P}\left(\frac{(u^{*})^{e(H)}}{|Aut(H)|}-\frac{p_{0}^{e(H)}}{|Aut(H)|}\right)^{2}$ (C.8)